OpenStudy (anonymous):

Explain how to graph x – 2y = 4 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only.

OpenStudy (anonymous):

do you know how to find the x and y intercept ?

OpenStudy (anonymous):

y= mx+b?

OpenStudy (anonymous):

to find the x intercept, sub in 0 for y. x - 2y = 4 x - 2(0) = 4 x - 0 = 4 x = 4 to find the y intercept, sub in 0 for x can you find the y intercept ?

OpenStudy (anonymous):

if you are looking for x let y=0 then if y let x=0

OpenStudy (anonymous):

then use ax+by=c

OpenStudy (anonymous):

i dont know the y-intercept but i know you sub 0

OpenStudy (anonymous):

x - 2y = 4 0 - 2y = 4 can you finish this to solve for y ?

OpenStudy (anonymous):

y = 2

OpenStudy (anonymous):

try again...don't forget about your negative sign

OpenStudy (anonymous):

y = -2

OpenStudy (anonymous):

the x-unit (4,0) the y unit (0,2)

OpenStudy (anonymous):

that is the graph

OpenStudy (anonymous):

correct....and as sammy said, x unit is (4,0) and y unit is (0,-2) plot those points and connect them with a line.

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

that is the graph

OpenStudy (anonymous):

but can you explain how to move on the graph

OpenStudy (anonymous):

you got a point on the x line on 4 and a point on the y line at -2. x - 2y = 4 change to y = mx + b -2y = -x + 4 (divide by -2) y = -1/-2x + 4/-2 y = 1/2x - 2 in y = mx + b, the m is your slope. Therefore your slope is 1/2 rise/run. now go to the -2 on your y axis. Go up 1 and to the right 2 and up 1 to the right 2 and you should reach the x axis at 4.

OpenStudy (anonymous):

does that make sense ?

OpenStudy (anonymous):

yes it does thanks a billion <3

OpenStudy (anonymous):

anytime :)